The local farmer's market offers one-kilogram asparagus for 28 €, two kilograms for 1,7 x 28 € and three kilograms for 1,9 x 281 €. What is the marginal cost of the second kilogram? What is the marginal cost of the third kilogram?


1
Expert's answer
2021-02-24T10:05:07-0500

Marginal cost represents the incremental costs incurred when producing additional units of a good or service. It is calculated by taking the total change in the cost of producing more goods and dividing that by the change in the number of goods produced.


"marginal\ cost \ =\frac{change \ in \ total\ cost}{change \ in \ quantity }"

cost of production at:

1 kilogram= "28"

2 kilogram="[1.7 \times 28]=47.6"

3 kilogram=["1.9\times 28]=53.2"


"marginal \ cost =\frac{change \ in \ total\ cost}{ change\ in\ quantity}"


at :

"2kgs =\frac{47.6-28}{2-1}=\frac{19.6}{1}=19.6\\

3kgs=\frac{53.2-47.6}{3-2}=\frac{5,6}{1}=5.6"